INAF queue observing, 2015 Jan 14-15

Observer: E. Sani, S. Bisogni
Telescope Operator: D. Gonzales Huerta
Support Astronomer: Michelle Edwards (on call)
Instrument: LUCI-MODS-LBC

Summary

During the twilight the dome is closed due to high humidity (>95%) and temperature clore to the dew point (-8.1C). Thin cirrus are passing, clouds on the NE horizon.
Once the dome is opened we have thin cirrus scattered over the sky.

Data Summary

Project_RUN Objects Filter ExpTime Notes
Antoniucci hip27881 (vytau standard) grat 200H+K filters HK - zJ 4s  
MIgnoli Feige34 red grat only filter s_SDSS 6min  
MIgnoli MOS22 red grat only filter s_SDSS 1.5 h  
MIgnoli LBC_red_MOS22 filter z_SDSS    

Details

06:32 the TO opens the dome.
  • Since we had some snow today, M. Wagner check for water/ice liking in the dome.

07:03 since the enclosure stay close several nights there are sting temperature gradients between the mirrors and insite/outside the dome. The TO cleans the optics

07:11 executeLUCIScript.sh hip27881_ls_zJ.acq luci.0002-0005 (standard for vytau target Antoniucci)

07:27 executeLUCIScript.sh hip27881_ls_zJ.spec luci.0006-0007

07:30 executeLUCIScript.sh hip27881_ls_HK.spec luci.0008-0009

07:34 the TO closes the dome because the temperature reaches the dew point and the humidity is 96%

09:08 Humidity drops down, the TO opens the dome again

09:45 executeLUCIScript.sh A1_LS_JHeI_tell.acq

09:56 TO turns on the pseudo monocular mode to allow Rome running tests for LBC collimation

09:58 LUCI mask got stuck; we contact the OSA.

10:20 We are not able to solve the problem with LUCI so we switch to MODS

10:36 slewing to Mignoli MOS22 standard star

10:48 acqMODS feige34.acq mods1r.0002-0003

11:00 execMODS feige34.obs mods1r.0004-0006

11:17 acqMODS mods.22..acq mods1r.0007-0009

11:34 execMODS mods.22.1.obs mods1r.0010-0013

11:43 we also start acquisition with LBC red (parallel mode)
  • collimation performed by LBC team in Rome
  • run 1_1_MOS22_z.ob
  • seeing 1.1''
  • 114324, 115108 show a jump due to offsetting with MODS

12:13 we stop the integration with LBC; it got stack at the fifth image

12:14 run 1_1_MOS22_z.ob
  • average seeing 1.1"
  • some images (111416 for example) seem to be affect by rotation effects due to to the fact the tracker doesn't work properly.
  • 12:39 average seeing 0.97''

12:45 execMODS mods.22.1.obs mods1r.0014-0015
  • 13:11 average seeing 1"

13:25 acqMODS Feige34.acq mods1r.0016 (mods_lastOffset)

13:32 execMODS feige34.obs mods1r.0017-0019 (last image 0019 to be discarded: only few seconds of integration)
  • average seeing 1.1"

13:57 running LBC flats; we are able to acquire only 1 frame in z band before the TO closes the dome

14:01 running LBC bias (25_Bias_Bino.ob)

-- %USERSIG{EleonoraSani - 2015-01-15}%

Comments

 
Topic revision: r10 - 19 May 2015, FeliceCusano
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback